Solve the square root of -2x+5=3


1
Expert's answer
2021-03-16T17:22:29-0400

"\sqrt{\smash[b]{-2x+5}}=3"

"(\sqrt{\smash[b]{-2x+5}})^2=3^2"

"-2x + 5 = 9"

"-2x = 9 -5=4"

"x = -4\/2 = -2"